Hello,

I need to find the slope of 3x+y=4.

Thanks

Answer by ptaylor(1894)

I need to find the slope of 3x+y=4.
This is the equation of a straight line. The standard form for this equation is :
y=mx+b
Where m is the slope and b is the y-intercept. This is called the slope- intercept form for a straight line.
Now we need to put your equation into standard form:
3x+y=4 subtract 3x from both sides
3x-3x+y=-3x+4
y=-3x+4 This is now in standard slope-intercept form and we can see that the slope is -3--------------ans

Hope this helps-----ptaylor